# LRT-Diffusion: Calibrated Risk-Aware Guidance for Diffusion Policies

### v1

@@ 3.3 Hypotheses for a Reverse Step
$H_{0}~(\text{background}):~~a_{t-1}\mid a_{t},s~\sim~\mathcal{N}\!\big(\mu_{u}(t,s,a_{t}),\,\sigma_{t}^{2}I\big),$
$H_{1}~(\text{good}):~~a_{t-1}\mid a_{t},s~\sim~\mathcal{N}\!\big(\mu_{c}(t,s,a_{t}),\,\sigma_{t}^{2}I\big),$
@@
3.4 Likelihood-Ratio View of a Reverse Step
Unknown labels $\Rightarrow$ a trajectory-level test.
At inference the label $c$ is unknown. Given two simple trajectory models $p_{u}(a_{T:1}\mid s)$ (background) and $p_{c}(a_{T:1}\mid s)$ (good), the Neyman–Pearson (NP) lemma implies that the level-$\alpha$ test that rejects $H_{0}$ when the cumulative log–likelihood ratio (LLR) exceeds a threshold is uniformly most powerful (UMP). We therefore base our gate on the trajectory statistic
@@ 3.4 Likelihood-Ratio View of a Reverse Step
How to compute $\ell_{\mathrm{cum}}$: step-wise LLRs.
With the reverse chain factorization in DDPM, for $i\in\{u,c\}$
@@
$\ell_{\mathrm{cum}}(a_{T:1})~=~\sum_{t=1}^{T}\ell_{t},\qquad\ell_{t}~:=~\log\frac{p_{c}(a_{t-1}\mid a_{t},s,t)}{p_{u}(a_{t-1}\mid a_{t},s,t)}.$ (4)
Head-induced one-step densities and quadratic form.
From Sec. 3.3, each head induces a Gaussian kernel with shared covariance. Hence the step-wise LLR admits the shared-variance simplification
$\ell_{t}~=~\frac{1}{2\sigma_{t}^{2}}\Big(\|a_{t-1}-\mu_{u}(t,s,a_{t})\|_{2}^{2}-\|a_{t-1}-\mu_{c}(t,s,a_{t})\|_{2}^{2}\Big),$ (5)
@@
Decision rule for assigning labels.
The NP test uses $\ell_{\mathrm{cum}}$ with a calibrated threshold $\tau$ to control the level-$\alpha$ Type-I rate. We implement a smooth gate that monotonically approximates the NP decision and calibrate $\tau$ under $H_{0}$; details are given in Sec. 4.
4 Method: LRT-Diffusion
@@
4.1 Good‑vs‑background Labeling
We rank state–action pairs by the IQL advantage $A(s,a)=\hat{Q}(s,a)-\hat{V}(s)$ computed on standardized inputs, and mark the top‑p global quantile as “good.” Although one could define state‑wise thresholds, the advantage already subtracts a state‑dependent baseline $\hat{V}(s)$, making A comparable across states in practice. Moreover, global top‑p is a non‑parametric surrogate of advantage‑weighted learning (e.g., AWR/AWAC), corresponding to a small‑temperature limit without tuning an extra temperature. Practically, offline datasets have highly uneven coverage—many states appear once or with only a handful of actions—so per‑state ranking/thresholding is statistically brittle (often ill‑posed); a single global threshold is more stable and reproducible. Crucially, our risk control via LRT is orthogonal to this choice: once the two‑head model is trained, Type‑I error is calibrated at inference regardless of how the conditional head’s subset was selected.
4.2 Two-head Training
@@
4.3 Evidence-gated Sampler
4.3.1 Motivation
1D evidence direction.
@@ 4.3.1 Motivation
From Eq. 5, under equal covariances the one-step LLR has constant gradient $\nabla_{a_{t-1}}\ell_{t}=\Sigma_{t}^{-1}(\mu_{c}-\mu_{u})$; among unit directions $u$ the directional derivative $\langle u,\nabla\ell_{t}\rangle$ is maximized by $u\parallel\Sigma_{t}^{-1}(\mu_{c}-\mu_{u})$ (isotropic: $u\parallel\mu_{c}-\mu_{u}$). We therefore restrict the reverse-step mean to the 1D ray $\mu_{u}+\beta(\mu_{c}-\mu_{u}),$ and let the scalar gate $\beta$ depend monotonically on the cumulative evidence $\ell_{\mathrm{cum}}$.
Why not always use $\mu_{c}$?
@@
Always setting $a_{t-1}{=}\mu_{c}{+}\sigma_{t}z_{t}$ is equivalent to always accepting $H_{1}$ (i.e., $\alpha{=}1$), which (i) removes any control on false activations of the conditional head and thus increases off-support mass in offline RL, and (ii) couples performance to critic/label errors where the model is most brittle. Our gate opens only when there is sufficient evidence (large $\ell_{\mathrm{cum}}$), yielding a calibrated budget on false activations (Prop. 2 and Thm. 2). Section 5.3 further shows that, when off-support errors dominate ($\varepsilon_{\rm out}\!\gg\!\varepsilon_{\rm in}$), reducing the state-conditional OOD rate via a smaller $\alpha$ tightens a lower bound on the true return (Prop. 3).
4.3.2 Gate Selection
Hard gate (UMP at level-$\alpha$).
@@ 4.3.2 Gate Selection
Let $\tau$ be a threshold. The hard likelihood-ratio test accepts $H_{1}$ when $\ell_{\mathrm{cum}}\geq\tau$ and rejects otherwise. Equivalently, with a cap $\beta_{\max}\!\in[0,1]$,
$\beta_{t}^{\text{h}}\;=\;\beta_{\max}\,\mathbf{1}\{\ell_{\mathrm{cum}}\geq\tau\},\quad a_{t-1}\;=\;\mu_{u}\,+\,\beta_{t}^{\text{h}}(\mu_{c}-\mu_{u})\,+\,\sigma_{t}z_{t}.$ (9)
@@
Under equal covariances and the joint factorization of the reverse chain, this hard test is uniformly most powerful (UMP) among all level-$\alpha$ tests (Prop. 1).
@@
Soft gate (stable surrogate) and its hard-limit.
For numerical stability we use a logistic surrogate
@@
$\beta_{t}\;=\;\beta_{\max}\,\sigma\!\Big(\tfrac{\ell_{\mathrm{cum}}-\tau}{\delta}\Big),\quad a_{t-1}\;=\;\mu_{u}\,+\,\beta_{t}(\mu_{c}-\mu_{u})\,+\,\sigma_{t}z_{t},$ (10)
@@
where $z_{t}\!\sim\!\mathcal{N}(0,I)$ and $\delta{>}0$ controls sharpness (smaller $\delta$ $\to$ sharper switch). As $\delta\!\downarrow\!0$, the soft rule converges to the hard gate both pointwise and in trajectory law (Lemma 1). Thus, the $\alpha,\tau,\beta_{t}$ can be seen as monotone, interpretable risk knobs. The resulting sampler is summarized in Alg. 2.
4.4 Calibration
Fix $(\beta_{\max},\delta)$ and freeze the sampler in Alg.2. On a held-out state set matched to deployment, simulate reverse chains under $H_{0}$ using the same sampler and collect the realized $\ell_{\mathrm{cum}}$; set
$\hat{\tau}=\mathrm{Quantile}_{1-\alpha}\big\{\ell_{\mathrm{cum}}^{(i)}\big\}.$
@@ 4.4 Calibration
Calibrating with the exact deployment sampler preserves the level-$\alpha$ semantics (Prop. 2). By the Dvoretzky–Kiefer–Wolfowitz bound Dvoretzky1956DKW ; Massart1990 , with probability $\geq 1-\eta$ over $n$ calibration draws,
@@
giving a finite-sample guarantee on the realized Type-I rate (Thm. 2). See Alg. 1 for the Monte Carlo procedure.
4.5 Inference
We calibrate $\tau$ once per (task, model, gate hyperparameters) on a held-out state set, using the same frozen sampler as in Alg. 2 (and Alg. 1 for the Monte-Carlo procedure). Thereafter, all rollouts simply call Alg. 2 with this fixed $\hat{\tau}$; no re-calibration is needed unless $(\beta_{\max},\delta)$ or the optional $Q$-composition change. The returned $a_{0}$ is on the standardized scale and is mapped to the environment as $a_{\mathrm{env}}=a_{0}\odot\mathrm{std}_{a}+\mathrm{mean}_{a}$ (then clipped to action bounds).
@@ 4.5 Inference
7:   Sample $s,a_{T}\sim\mathcal{N}(0,I)$; $\ell\leftarrow 0$
@@
9:     $(\mu_{u},\mu_{c})\leftarrow\mu_{\theta}(a_{t},s)$
10:     $\beta_{t}\leftarrow\beta_{\max}\sigma((\ell-\tau)/\delta)$(soft-gate)
@@
5:  $(\mu_{u},\mu_{c})\leftarrow\mu_{\theta}(a_{t},s)$
6:  $\beta_{t}\leftarrow\beta_{\max}\sigma((\ell-\tau)/\delta)$
@@
4.6 Composition with Value Gradients
To cleanly separate risk control (via the LRT gate) from return seeking, we optionally add a small critic step that mirrors common practice in diffusion-policy samplers. This step uses the action-space gradient of a learned critic (not the diffusion score) with a simple schedule and clipping:
@@ 4.6 Composition with Value Gradients
We use the evaluation center $a_{c}=\mu_{\text{LRT},t}$ as evidence opens the gate, suggesting the gradient is closer to the good direction. For more evaluation center choice discussion, refer Appendix D.1
5 Theory and Properties
5.1 Gate Selection: UMP vs. Stability
In this part, we discuss the properties of hard and soft gates. For the full mathematical proof, see Appendix A.
Proposition 1 (Neyman–Pearson optimality).
@@ 5 Theory and Properties
Proposition 2 (Calibrated semantics under the deployment sampler).
@@
up to finite-sample fluctuations (see Thm. 2 for a DKW bound).
Discussion.
The hard gate is UMP at level $\alpha$ (Prop. 1); the soft gate (Lemma 1) provides a numerically stable approximation with the same level-$\alpha$ semantics when $\tau$ is calibrated under the exact deployment sampler (Prop. 2).
5.2 Finite-sample Calibration and Stability
@@
5.2.1 Finite-sample Guarantee
Theorem 2 (Calibration accuracy via DKW Dvoretzky1956DKW ; Massart1990 ).
For any $\zeta\in(0,1)$, with probability at least $1-\zeta$ over the calibration sample,
$\mathbb{P}_{H_{0}}^{\text{(sampler)}}\!\big(\ell_{\mathrm{cum}}\geq\hat{\tau}\big)~\leq~\alpha+\varepsilon_{n},\qquad\varepsilon_{n}=\sqrt{\tfrac{1}{2n}\log\tfrac{2}{\zeta}}.$
Calibration budget. Rearranging the bound gives a one-line rule: to guarantee $\mathbb{P}_{H_{0}}^{(\text{sampler})}(\ell_{\mathrm{cum}}\!\geq\!\hat{\tau})\leq\alpha+\varepsilon$ with confidence at least $1-\zeta$, it suffices to set sample size $n\;\geq\;\frac{1}{2\varepsilon^{2}}\,\log\!\frac{2}{\zeta}.$ See Appendix B.1 for proof.
5.2.2 Stability
Lemma 3 (Deterministic displacement bound).
@@ 5.2.2 Stability
$\|m_{t}-\mu_{u,t}\|~\leq~\beta_{\max}\,\|\Delta\mu_{t}\|\;+\;\lambda_{\max}\,\sigma_{t}^{2}\,G\;=:\;B_{t}.$ (12)
@@
Thus, the deterministic component of the sampler admits a uniform, hyperparameter‑controlled drift bound both per step and over the full reverse trajectory; see Appendix B.2 for the detailed proof.
Implication for one-step LLR fluctuations.
Under equal covariances, the one-step LLR increment admits the identity
$\Delta\ell_{t}~=$~ $\underbrace{\frac{\Delta\mu_{t}^{\top}\big(m_{t}-\tfrac{\mu_{c,t}+\mu_{u,t}}{2}\big)}{\sigma_{t}^{2}}}_{\text{deterministic part}}\;+\;\underbrace{\frac{\Delta\mu_{t}^{\top}z_{t}}{\sigma_{t}}}_{\text{zero-mean Gaussian}},$
@@
$\big|\mathbb{E}[\Delta\ell_{t}]\big|$ $~\leq~\frac{\|\Delta\mu_{t}\|}{\sigma_{t}^{2}}\,\Big(B_{t}+\tfrac{1}{2}\|\Delta\mu_{t}\|\Big),$
$\mathrm{Var}(\Delta\ell_{t})$ $=~\frac{\|\Delta\mu_{t}\|^{2}}{\sigma_{t}^{2}}.$
In particular, if $\|\Delta\mu_{t}\|\leq D$ and $\sigma_{t}\geq\sigma_{\min}>0$, then $\mathrm{Var}(\Delta\ell_{t})\leq D^{2}/\sigma_{\min}^{2}$ and $\big|\mathbb{E}[\Delta\ell_{t}]\big|\leq\frac{D}{\sigma_{\min}^{2}}\big(B_{\text{step}}+\tfrac{D}{2}\big)$, so $\ell_{\mathrm{cum}}=\sum_{t}\Delta\ell_{t}$ is sub-Gaussian with variance proxy $\sum_{t}\|\Delta\mu_{t}\|^{2}/\sigma_{t}^{2}\leq T\,D^{2}/\sigma_{\min}^{2}$. This quantifies that our gate and gradient clipping keep both the magnitude of mean shifts and the variance of the accumulated evidence controlled. See proofs on Appendix B.3 and  B.4.
5.3 Distributional Control and Return Bounds

### v2

@@ 3.3. Hypotheses for a Reverse Step
$H_{0}\penalty 10000\ (\text{background}):\penalty 10000\ \penalty 10000\ a_{t-1}\mid a_{t},s\penalty 10000\ \sim\penalty 10000\ \mathcal{N}\!\big(\mu_{u}(t,s,a_{t}),\,\sigma_{t}^{2}I\big),$
$H_{1}\penalty 10000\ (\text{good}):\penalty 10000\ \penalty 10000\ a_{t-1}\mid a_{t},s\penalty 10000\ \sim\penalty 10000\ \mathcal{N}\!\big(\mu_{c}(t,s,a_{t}),\,\sigma_{t}^{2}I\big),$
@@
3.4. Likelihood-Ratio View of a Reverse Step
Unknown labels $\Rightarrow$ a trajectory-level test. At inference the label $c$ is unknown. Given two trajectory models $p_{u}(a_{T:1}\mid s)$ (background) and $p_{c}(a_{T:1}\mid s)$ (good), the Neyman–Pearson (NP) lemma implies that the level-$\alpha$ test that rejects $H_{0}$ when the cumulative log–likelihood ratio (LLR) exceeds a threshold is uniformly most powerful (UMP). We thus base our gate on the trajectory statistic
@@ 3.4. Likelihood-Ratio View of a Reverse Step
How to compute $\ell_{\mathrm{cum}}$: step-wise LLRs. With the reverse chain factorization in DDPM, for $i\in\{u,c\}$
@@
$\ell_{\mathrm{cum}}(a_{T:1})\penalty 10000\ =\penalty 10000\ \sum_{t=1}^{T}\ell_{t},\qquad\ell_{t}\penalty 10000\ :=\penalty 10000\ \log\frac{p_{c}(a_{t-1}\mid a_{t},s,t)}{p_{u}(a_{t-1}\mid a_{t},s,t)}.$ (4)
Head-induced one-step densities and quadratic form. From Sec. 3.3, each head induces a Gaussian kernel with shared covariance. Hence the step-wise LLR admits the shared-variance simplification
$\ell_{t}\penalty 10000\ =\penalty 10000\ \frac{1}{2\sigma_{t}^{2}}\Big(\|a_{t-1}-\mu_{u}(t,s,a_{t})\|_{2}^{2}-\|a_{t-1}-\mu_{c}(t,s,a_{t})\|_{2}^{2}\Big),$ (5)
@@
Decision rule for assigning labels. The NP test uses $\ell_{\mathrm{cum}}$ with a calibrated threshold $\tau$ to control the level-$\alpha$ Type-I rate. We implement a smooth gate that monotonically approximates the NP decision and calibrate $\tau$ under $H_{0}$; details are given in Sec. 4.
4. Method: LRT-Diffusion
@@
4.1. Good‑vs‑background Labeling
We rank state–action pairs by the IQL advantage $A(s,a)=\hat{Q}(s,a)-\hat{V}(s)$ computed on standardized inputs, and mark the top‑p global quantile as “good.” Although one could define state‑wise thresholds, the advantage already subtracts a state‑dependent baseline $\hat{V}(s)$, making A comparable across states in practice. Moreover, global top‑p is a non‑parametric surrogate of advantage‑weighted learning (e.g., AWR/AWAC), corresponding to a small‑temperature limit without tuning an extra temperature. Practically, offline datasets have highly uneven coverage—many states appear once or with only a handful of actions—so per‑state ranking is statistically brittle; a single global threshold is more stable and reproducible. Crucially, our risk control via LRT is orthogonal to this choice: once the two‑head model is trained, Type‑I error is calibrated at inference regardless of how the conditional head’s subset was selected.
4.2. Two-head Training
@@
4.3. Evidence-gated Sampler
4.3.1. Motivation
@@
1D evidence direction. From Eq. 5, under equal covariances the one-step LLR has constant gradient $\nabla_{a_{t-1}}\ell_{t}=\Sigma_{t}^{-1}(\mu_{c}-\mu_{u})$; among unit directions $u$ the directional derivative $\langle u,\nabla\ell_{t}\rangle$ is maximized by $u\parallel\Sigma_{t}^{-1}(\mu_{c}-\mu_{u})$ (isotropic: $u\parallel\mu_{c}-\mu_{u}$). We therefore restrict the reverse-step mean to the 1D ray $\mu_{u}+\beta(\mu_{c}-\mu_{u}),$ and let the scalar gate $\beta$ depend monotonically on the cumulative evidence $\ell_{\mathrm{cum}}$.
@@
Why not always use $\mu_{c}$? Always setting $a_{t-1}{=}\mu_{c}{+}\sigma_{t}z_{t}$ is equivalent to always accepting $H_{1}$, which (i) removes any control on false activations of the conditional head and thus increases off-support mass in offline RL, and (ii) couples performance to critic/label errors where the model is most brittle. Our gate opens only when there is sufficient evidence (large $\ell_{\mathrm{cum}}$), yielding a calibrated budget on false activations (Prop. 5.1 and Thm. 5.2.1). Section 5.3 further shows that, when off-support errors dominate ($\varepsilon_{\rm out}\!\gg\!\varepsilon_{\rm in}$), reducing the state-conditional OOD rate via a smaller $\alpha$ tightens a lower bound on the true return (Prop. 5.3).
4.3.2. Gate Selection
@@
Hard gate (UMP at level-$\alpha$). Let $\tau$ be a threshold. The hard likelihood-ratio test accepts $H_{1}$ when $\ell_{\mathrm{cum}}\geq\tau$ and rejects otherwise. Equivalently, with a cap $\beta_{\max}\!\in[0,1]$,
$\beta_{t}^{\text{h}}\;=\;\beta_{\max}\,\mathbf{1}\{\ell_{\mathrm{cum}}\geq\tau\},\quad a_{t-1}\;=\;\mu_{u}\,+\,\beta_{t}^{\text{h}}(\mu_{c}-\mu_{u})\,+\,\sigma_{t}z_{t}.$
@@
Under equal covariances and the joint factorization of the reverse chain, this hard test is uniformly most powerful (UMP) among all level-$\alpha$ tests (Prop. 5.1).
@@
Soft gate (stable surrogate) and its hard-limit. For numerical stability we use a logistic surrogate
@@
$\beta_{t}\;=\;\beta_{\max}\,\text{Sigmoid}\!\left(\!\tfrac{\ell_{\mathrm{cum}}-\tau}{\delta}\right),\quad a_{t-1}\;=\;\mu_{u}\,+\,\beta_{t}(\mu_{c}-\mu_{u})\,+\,\sigma_{t}z_{t},$
@@
where $z_{t}\!\sim\!\mathcal{N}(0,I)$ and $\delta{>}0$ controls sharpness (smaller $\delta$ $\to$ sharper switch). As $\delta\!\downarrow\!0$, the soft rule converges to the hard gate both pointwise and in trajectory law (Lemma 5.1). Thus, the $\alpha,\tau,\beta_{t}$ can be seen as monotone, interpretable risk knobs. The resulting sampler is summarized in Alg. 2.
4.4. Calibration
@@
Fix $(\beta_{\max},\delta)$ and freeze the sampler in Alg.2. On a held-out state set matched to deployment, simulate reverse chains under $H_{0}$ using the same sampler and collect the realized $\ell_{\mathrm{cum}}$; set $\hat{\tau}=\mathrm{Quantile}_{1-\alpha}\big\{\ell_{\mathrm{cum}}^{(i)}\big\}$. Calibrating with the exact deployment sampler preserves the level-$\alpha$ semantics (Prop. 5.1). By the Dvoretzky–Kiefer–Wolfowitz bound Dvoretzky1956DKW; Massart1990, with probability $\geq 1-\eta$ over $n$ calibration draws,
@@
giving a finite-sample guarantee (Thm. 5.2.1); see Alg. 1.
4.5. Inference
We calibrate $\tau$ once per (task, model, gate hyperparameters) on a held-out state set, using the same frozen sampler as in Alg. 2 (and Alg. 1 for the Monte-Carlo procedure). Thereafter, all rollouts simply call Alg. 2 with this fixed $\hat{\tau}$; no re-calibration is needed unless hyperparameters change. The final $a_{0}$ is un-standardized and clipped before deployment (Alg. 2).
@@ 4.5. Inference
7:   Sample $s\sim\mathrm{Unif}(\mathcal{S}_{\rm cal}),a_{T}\sim\mathcal{N}(0,I)$; $\ell\leftarrow 0$
@@
9:     $(\mu_{u},\mu_{c})\leftarrow\mu_{\theta}(t,s,a_{t})$
10:     $\beta_{t}\leftarrow\beta_{\max}\text{Sigmoid}\!\left((\ell-\tau)/\delta\right)$(soft-gate)
@@
5:  $(\mu_{u},\mu_{c})\leftarrow\mu_{\theta}(t,s,a_{t})$
6:  $\beta_{t}\leftarrow\beta_{\max}\text{Sigmoid}\!\left((\ell-\tau)/\delta\right)$
@@
4.6. Composition with Value Gradients
To cleanly separate risk control from return seeking, we optionally apply a value-guidance step. Specifically, we ascend the action-space gradient of a learned critic with a simple schedule and clipping:
@@ 4.6. Composition with Value Gradients
We use the evaluation center $a_{c}=\mu_{\text{LRT},t}$ as evidence opens the gate, suggesting the gradient is closer to the good direction. For more evaluation center choice discussion, refer Appendix E.1
5. Theory and Properties
5.1. Gate Selection: UMP vs. Stability
In this part, we discuss the properties of hard and soft gates. For the full mathematical proof, see Appendix B.
[…]
@@ 5. Theory and Properties
[…]
@@
up to finite-sample fluctuations (see Thm. 5.2.1 for a DKW bound).
5.2. Finite-sample Calibration and Stability
@@
5.2.1. Finite-sample Guarantee
{theorem}
[Calibration accuracy via DKW Dvoretzky1956DKW; Massart1990] For any $\zeta\in(0,1)$, with probability at least $1-\zeta$ over the calibration sample,
$\mathbb{P}_{H_{0}}^{\text{(sampler)}}\!\big(\ell_{\mathrm{cum}}\geq\hat{\tau}\big)\penalty 10000\ \leq\penalty 10000\ \alpha+\varepsilon_{n},\qquad\varepsilon_{n}=\sqrt{\tfrac{1}{2n}\log\tfrac{2}{\zeta}}.$
Rearranging the bound gives a one-line rule: to guarantee
$\mathbb{P}_{H_{0}}^{(\text{sampler})}(\ell_{\mathrm{cum}}\!\geq\!\hat{\tau})\leq\alpha+\varepsilon$ with confidence at least $1-\zeta$, it suffices to set sample size $n\;\geq\;\frac{1}{2\varepsilon^{2}}\,\log\!\frac{2}{\zeta}.$ See Appendix C.1 for proof.
5.2.2. Stability
[…]
@@ 5.2.2. Stability
$\|m_{t}-\mu_{u,t}\|\penalty 10000\ \leq\penalty 10000\ \beta_{\max}\,\|\Delta\mu_{t}\|\;+\;\lambda_{\max}\,\sigma_{t}^{2}\,G\;=:\;B_{t}.$ (10)
@@
Thus, the deterministic component of the sampler admits a uniform, hyperparameter‑controlled drift bound both per step and over the full reverse trajectory; see Appendix C.2 for the detailed proof.
Implication for one-step LLR fluctuations. Under equal covariances, the one-step LLR increment admits the identity
$\Delta\ell_{t}\penalty 10000\ =\penalty 10000\$$\underbrace{\frac{\Delta\mu_{t}^{\top}\big(m_{t}-\tfrac{\mu_{c,t}+\mu_{u,t}}{2}\big)}{\sigma_{t}^{2}}}_{\text{deterministic part}}\;+\;\underbrace{\frac{\Delta\mu_{t}^{\top}z_{t}}{\sigma_{t}}}_{\text{zero-mean Gaussian}},$
@@
$\big|\mathbb{E}[\Delta\ell_{t}]\big|$ $\penalty 10000\ \leq\penalty 10000\ \frac{\|\Delta\mu_{t}\|}{\sigma_{t}^{2}}\,\Big(B_{t}+\tfrac{1}{2}\|\Delta\mu_{t}\|\Big),$
$\mathrm{Var}(\Delta\ell_{t})$ $=\penalty 10000\ \frac{\|\Delta\mu_{t}\|^{2}}{\sigma_{t}^{2}}.$
In particular, if $\|\Delta\mu_{t}\|\leq D$ and is the fixed schedule $\sigma_{t}\geq\sigma_{\min}>0$, then $\mathrm{Var}(\Delta\ell_{t})\leq D^{2}/\sigma_{\min}^{2}$ and $\big|\mathbb{E}[\Delta\ell_{t}]\big|\leq\frac{D}{\sigma_{\min}^{2}}\big(B_{\text{step}}+\tfrac{D}{2}\big)$, so $\ell_{\mathrm{cum}}=\sum_{t}\Delta\ell_{t}$ is sub-Gaussian with variance proxy $\sum_{t}\|\Delta\mu_{t}\|^{2}/\sigma_{t}^{2}\leq T\,D^{2}/\sigma_{\min}^{2}$. This quantifies that our gate and gradient clipping keep both the magnitude of mean shifts and the variance of the accumulated evidence controlled. See proofs on Appendix C.3 and C.4.
5.3. Distributional Control and Return Bounds